OpenStudy (anonymous):

solve the differential equation \[\frac{dy}{d\theta}=\frac{e^ysin^2\theta}{ysec\theta}\] \[\int\frac{y}{e^y}dy=\int\frac{sin^2\theta}{sec\theta}d\theta\] \[-e^{-y}(y+1)=\frac{sin^3\theta}{3}+C\] Where from here?

OpenStudy (lgbasallote):

isnt that the answer already o.O

OpenStudy (anonymous):

Hey Lgba!!!!! No I have to solve for y

OpenStudy (lgbasallote):

you have? well that's a bummer

OpenStudy (anonymous):

LOl, that's what I thought the e^y makes it ugly

OpenStudy (lgbasallote):

\[\frac{y+1}{e^y} = -\frac{\sin^3 \theta}{3} + C\] im not doing anything yet..just looking at a diff perspective

OpenStudy (anonymous):

Sure

OpenStudy (lgbasallote):

well you can ln it..

OpenStudy (anonymous):

\[ln|-y(y+1)|\]?

OpenStudy (anonymous):

\[yln|y+1|\]

OpenStudy (lgbasallote):

wait... this thing is \[\Large \int \frac{\sin^3 \theta}{\sec theta} d\theta\] right?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

should be \[\frac{\sin^4 \theta}{4} + C\] then

OpenStudy (anonymous):

well according to wolfram it's (sin^3)/3

OpenStudy (lgbasallote):

oh wait it's a two

OpenStudy (lgbasallote):

nevermind

OpenStudy (lgbasallote):

well i officially have no idea lol.

OpenStudy (anonymous):

yes you do....let's work on it a little more

OpenStudy (anonymous):

would \[\frac1y|y+1|\] be correct for the left side

OpenStudy (lgbasallote):

even when in ln it makes no sense \[y + \ln (y+1) = \ln (-\frac{\sin^3 \theta}{3} + C)\]

OpenStudy (lgbasallote):

\[\ln (e^y (y+1) ) \implies \ln e^y + \ln (y+1)\]

OpenStudy (anonymous):

Did I make an error in the initial integration?

OpenStudy (lgbasallote):

should be -y

OpenStudy (anonymous):

\[\int \frac{y}{e^y}dy\]=?

OpenStudy (lgbasallote):

your integral is right

OpenStudy (anonymous):

\[y=-1-W \left (\frac{1}{3e} \sin^3 (\theta ) +C\right )\] Where W(z) is Lambert's W function, the inverse function of z(w)=we^w. You can do it, but it requires a special function not typically used in lower-level maths. You did your integration right, so you probably just wrote the problem wrong. Also, it is common practice in diff eqns to leave solutions in their implicit form, so I doubt that your teacher really wanted you to solve for y.

OpenStudy (lgbasallote):

exactly :| leave in implicit form lol

OpenStudy (anonymous):

Let's start from the beginning \[\frac{dy}{d\theta}=\frac{e^ysin^2\theta}{ysec\theta}\]

OpenStudy (anonymous):

\[-e^{-y}(y+1)=\frac{sin^3\theta}{3}+C\] I'll leave this as the final answer.